Your investment will pay you $15,000 at the end of one year, $14,000 at the end of the following year, and $24,000 at the end of the year after that (three years from today). The interest rate is 4%. (a) What is the present value? (b) What will be the future value three years later ? (c) Assuming that the investment costs $40,000 today, what is the net present value (NPV) of the investment?
Answer #1

Present Value = 15,000/(1.04) + 14,000/(1.04)2 + 24,000/(1.04)3

Present Value = $48,702.78

Future Value = 15,000(1.04)2 + 14,000(1.04) + 24,000

Future Value = $54,784

NPV = -40,000 + 48,702.78

NPV = 8,702.78